OpenStudy (anonymous):

Please help for test!!!! A circle with center at (-2,3) and the circle is tangent to the line x=5

OpenStudy (anonymous):

I have the (x+2)^2 + (y-3)^2=? part but I don't know what to do with the x=5 part

OpenStudy (anonymous):

derivative of the circle function gives tangent line plug in 5 into the derivative and you get y value what's your question exactly?

OpenStudy (anonymous):

i think..idunno not too familiar with circles

OpenStudy (anonymous):

How do I get the part with the "?" using the tangent line info. This is precalc so no derivatives please

OpenStudy (anonymous):

Remember that a tangent line means one point on the circle is touches the line x=5. Since it is a circle, all the points are an equal distance away from the center.  the distance of the line from the center is 5 - (-2) = 7 so the radius is 7 which means (x+2)^2 + (y+3)^2 = 7^2 (x+2)^2 + (y+3)^2 = 49